OpenStudy (anonymous):

Consider the line which passes through the point P(-3, 5, 1), and which is parallel to the line x=1+6t,y=2+5t,z=3+4t Find the point of intersection of this new line with each of the coordinate planes: I've got <1,2,3> + t<6,5,4> so far, but not sure where really to go from there.

OpenStudy (anonymous):

someone???? anyone?????? seriously?????? due in like a few hours................and this is only question 5 of 20.........................

OpenStudy (anonymous):

tried moving on, but now i'm also stuck on another problem because these stupid questions aren't explained in the stupid book that i paid 200 stupid bucks for!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (rational):

where are you stuck

OpenStudy (anonymous):

i said where i was stuck......said this is what i have, don't know where to go from there.

OpenStudy (anonymous):

@True_Alpha_Wolf @Jman00000 @surd @Punkchick @chainedecho

OpenStudy (rational):

the required line is <-3, 5, 1> + t<6,5,4> right ?

OpenStudy (true_alpha_wolf):

im bad at math

OpenStudy (anonymous):

ready to cry or stab myself in the eye with my pencil here. first i deal with someone cursing me out for no reason, then nothing at all

OpenStudy (anonymous):

y u mention me dafuq

OpenStudy (anonymous):

I sorry thought she could use the help

OpenStudy (rational):

you did not reply me properly in previous thread

OpenStudy (anonymous):

no, i had <1,2,3> + t<6,5,4>

OpenStudy (anonymous):

i'm sorry, @rational, i wasn't gonna sift through that garbage that other person was posting. i just ignored it and posted again

OpenStudy (rational):

thats the given line

OpenStudy (anonymous):

i'm already having a bad day and if i'd read the obscenities that guy was throwing around, i would have lost it.

OpenStudy (rational):

a line parallel to this will be having same direction vector, yes ?

OpenStudy (rational):

you're free to ignore him

OpenStudy (anonymous):

bye bye

OpenStudy (anonymous):

Tourists who had paid money for a safari were packed into the bus like sardines in a can. Which statement is true? The sentence features figurative language in the form of a metaphor. The sentence features figurative language in the form of a simile. The sentence features literal language in the form of a metaphor. The sentence features literal language in the form of simile.

OpenStudy (rational):

we're talking about some spammer who bugged us in the previous thread @sambolin1996

OpenStudy (anonymous):

oh lol

OpenStudy (rational):

anyways, did you get the direction vector concept ? @saiken2009

OpenStudy (anonymous):

i don't get it, why is my thing wrong? i did it just how the book told me to do it.

OpenStudy (anonymous):

Hey I was just trying to help :(

OpenStudy (rational):

all parallel lines will have same direction vector

OpenStudy (anonymous):

and the x0+at is there

OpenStudy (anonymous):

x0+at, etc, <x0,y0,z0>+t<a,b,c>

OpenStudy (rational):

given line : <1,2,3> + t<6,5,4> a line parallel to this line will be of form <x0, y0, z0> + t<6,5,4> yes?

OpenStudy (anonymous):

so if we have 1+6t=x, x0 should be 1 and a should be 6

OpenStudy (rational):

you're right

OpenStudy (rational):

see if u can answer my q above

OpenStudy (anonymous):

umm....

OpenStudy (anonymous):

ok, so yeah. i guess.

OpenStudy (anonymous):

little lost on this whole section. sorry if i'm a little slow.

OpenStudy (rational):

<x0, y0, z0> + t<6,5,4> since you know that the line passes throughP(-3, 5, 1), the equation would be : <-3, 5, 1> + t<6,5,4>

OpenStudy (rational):

its okay..

OpenStudy (anonymous):

so we just change the point of origin?

OpenStudy (rational):

yes, parallel lines will have same direction. just the point changes

OpenStudy (anonymous):

it's another line and the second part of that (the a, b, c) is just the vector that says the direction and mag and the point in the beginning (the x0, y0, z0) gives us a point in the space for the vector

OpenStudy (anonymous):

am i understanding that right?

OpenStudy (rational):

yes : a line can be represented using a "point" and a "direction vector" 

OpenStudy (rational):

for parallel lines, direction vector will be same only the point changes

OpenStudy (anonymous):

so the line i'm actually focusing on is <-3, 5, 1> +t<6,5,4>

OpenStudy (rational):

yes <-3, 5, 1> + t<6,5,4> and any point on this line will be : (-3+6t, 5+5t, 1+4t)

OpenStudy (anonymous):

and i know the xy plane coordinate will be (_, _, 0) and the xz will be (_,0,_).

OpenStudy (rational):

Correct ! for the intersection point with xy plane, simply set your z coordiante to 0 and solve t

OpenStudy (rational):

(-3+6t, 5+5t, `1+4t`)

OpenStudy (rational):

set that equal to 0 and solve t

OpenStudy (anonymous):

that gives me t though. how does that give me the x, y points in the plane?

OpenStudy (rational):

plug that t in the point again : (-3+6t, 5+5t, 1+4t)

OpenStudy (anonymous):

oh, so plug in -4 for the other two then.

OpenStudy (rational):

1+4t = 0 t = ?

OpenStudy (anonymous):

and so on for the other planes? but with different zeros

OpenStudy (anonymous):

ok, this is nowhere in my book and my teacher didn't do this. so frustrating.

OpenStudy (rational):

i knw

OpenStudy (anonymous):

and since i have a headache, that should have been -1/4 instead of t=4

OpenStudy (anonymous):

oops

OpenStudy (rational):

(-3+6t, 5+5t, 1+4t) you should get below intersecting points : xy plane : (-9/2, 15/4, 0) yz plane : (0, 15/2, 3) zx plane : (-9, 0, -3)

OpenStudy (anonymous):

thanks. @rational